OpenStudy (anonymous):

Which of the following points lies on the line whose equation is y = 2x - 3? (0, 3)??

OpenStudy (anonymous):

3 = 2(0) - 3 3 = 0 - 3 3 = -3 NO, so the point is not on the line.

OpenStudy (anonymous):

Are there any other options?